OpenStudy (anonymous):

What is the simplified form of 5x plus 4 over x plus 7minus the fraction 2x plus 1 over x plus 7 ? a)3x plus 3 over x plus 7 b)3x plus 5 over x plus 7 c)7x plus 3 over x plus 7 d)7x plus 5 over x plus 7 @Mathematics

OpenStudy (kropot72):

\[\frac{5x+4}{x+7}-\frac{2x+1}{x+7}\] x + 7 is a common denominator.\[\frac{5x+4-(2x+1)}{x+7}=\frac{5x+4-2x-1}{x+7}\] Can you finish it now?

OpenStudy (anonymous):

I got 3x+5 over x+7?

OpenStudy (kropot72):

\[\frac{5x+4-2x-1}{x+7}=\frac{3x+3}{x+7}\]

OpenStudy (anonymous):

Oh wow I feel silly now... I made a simple error haha. Thank you!

OpenStudy (kropot72):

You're welcome :)